OpenStudy (anonymous):

one equation in a linear system is 5x+3y=14. write another equation so that the system has the unique solution (4,-2)

OpenStudy (asnaseer):

simplest would be: x + y = 2

OpenStudy (asnaseer):

sorry - I mean obtained by just adding the x and y values of the SINGLE solution provided

OpenStudy (asnaseer):

does it make sense?

OpenStudy (anonymous):

i gues.. im not sure

OpenStudy (asnaseer):

you are told that the unique solution is: x=4, y=-2 so we know x + y = 4 + (-2) = 4 - 2 = 2

OpenStudy (asnaseer):

we could have created any other combination - as long as it is NOT a multiple of the 1st solution given - otherwise you would not be able to solve the system of simultaneous equations

OpenStudy (asnaseer):

e.g. 10x + 6y = 10*4 + 6*(-2) = 40 - 12 = 28 but this would not allow you to solve the system to find x and y.

OpenStudy (anonymous):

ahaa

OpenStudy (anonymous):

so that would be it ?>

OpenStudy (asnaseer):

good - I'm glad I was able to explain it clearly enough for you :-)

OpenStudy (anonymous):

ok so the solution would be x+y=4+(-2)

OpenStudy (asnaseer):

not the "solution", but the answer to the question - yes

OpenStudy (anonymous):

aaa ok yea i meant that

OpenStudy (asnaseer):

each equation represents a straight line. if the two lines are NOT parallel, then they will intersect. the point at which they intersect will be the "solution" to the system of equations.

OpenStudy (anonymous):

thank you i appreciate

OpenStudy (asnaseer):

yw